### v2

# Automatically Benchmarking LLM Code Agents through Agent-driven Annotation and Evaluation

Abstract.
Recent advances in code agents have enabled automated software development at the project level, supported by large language models (LLMs). However, existing benchmarks for code agent evaluation face two major limitations. First, creating high-quality project-level evaluation datasets requires extensive domain expertise, leading to prohibitive annotation costs and limited diversity. Second, while recent Agent-as-a-Judge paradigms address the rigidity of traditional unit tests by enabling flexible metrics, their reliance on In-Context Learning (ICL) with general LLMs often results in inaccurate assessments that misalign with human standards. To address these challenges, we propose an agent-driven benchmark construction pipeline that leverages human supervision to efficiently generate diverse project-level tasks. Based on this, we introduce PRDBench, comprising 50 real-world Python projects across 20 domains, each with structured Product Requirement Documents (PRDs) and comprehensive criteria. Furthermore, to overcome the inaccuracy of general LLM judges, we propose a highly reliable evaluation framework powered by a specialized, fine-tuned model. Based on Qwen3-Coder-30B, our dedicated PRDJudge achieves over 90% human alignment in fixed-interface scenarios. Extensive experiments demonstrate that our suite provides a scalable, robust, and highly accurate framework for assessing state-of-the-art code agents.
Key words and phrases:
Code Agent, Evaluation, Agent-as-a-Judge
doi:
HJFB4234
1. Introduction
In recent years, code agents have made significant progress, capable of solving increasingly complex programming tasks. From initially focusing on single-file code generation to now advancing towards complete project-level software development, Code Agents are demonstrating powerful capabilities in building comprehensive systems. LLM-based Code Agents such as CursorAgent anysphere2025cursoragent, Claude Code anthropic2025claude37 and Gemini CLI mullen2025geminicli have been widely adopted in real-world development scenarios, accelerating the proliferation of automated programming tools.
As code agents continue to advance, a variety of benchmarks have emerged to evaluate their capabilities. Several works chan2024mlebench; huang2024dacodeagentdatascience; daghighfarsoodeh2025deepbenchdeeplearningbenchmark focus on assessing LLMs and code agents in data science tasks, including subtasks such as data processing and machine learning training. Other benchmarks jimenez2023swe; fu2025corecodebench; vergopoulos2025automatedbenchmarkgenerationrepositorylevel select complex code repositories from GitHub pull requests and transform them, often with minimal human intervention, into testable scenarios. In addition, PaperBench starace2025paperbench evaluates the ability of agents to reproduce paper-level code implementations, providing a more rigorous assessment of their research reproducibility.
Benchmark Area # Project Annotator # Metrics Judger Claude Code Score
SWE-Bench jimenez2023swe Pull Requests 12 Web Crawler 2,294 Unit Test 70.3%
MLEBench chan2024mlebench AI Competition 75 Human 75 Test Set 51.1%
DevAI zhuge2024agent Developement 55 Human 365 Agent 73.0%
PaperBench starace2025paperbench AI Research 20 Human (PhD.) 8,316 Human/LLM 21.0%
PRDBench Engineering Development 50 Agent & Human 1,262 Agent (PRDJudge) 45.5%
Table 1. Comparison between PRDBench and project-level code agent benchmarks. We employ the performance of code agents with Claude as a benchmark indicator for assessing the task difficulty of our proposed evaluation suite. The results from external report anthropic2025claude37 is denoted by *, from our implementation is denoted by $\dagger$, otherwise we use the best results from the original paper with author specified Claude-driven agents.
Despite these advancements, current benchmarks face two major limitations. First, the creation of high-quality evaluation datasets requires substantial human effort and expertise. Generating reliable unit tests or grading tasks often depends on expert annotators, especially when migrating real-world scenarios such as Kaggle competitions or GitHub pull requests. These tasks demand domain-level verification and annotation to produce usable test cases. As code agents advance rapidly, the requirements for annotators’ expertise and time investment continue to increase. For example, PaperBench starace2025paperbench recruited ICML authors as annotators, with each annotation task demanding several days of work from PhD-level experts. This process not only incurs high annotation costs, but also restricts the diversity of evaluation data, as recruiting experts from various domains is challenging and often leads to datasets being sourced from a single domain.
Second, the rigidity of the evaluation limits the broad applicability of current benchmarks. Existing benchmarks jimenez2023swe primarily rely on unit test pass rates to assess code agent performance. While unit tests are effective for verifying specific functions or components, they are limited in scope and overly restrictive, enforcing strict requirements on project interfaces and implementation details. To address this rigidity and introduce broader validation methods, recent works have explored LLM-as-a-Judge or Agent-as-a-Judge paradigms. However, these approaches predominantly rely on In-Context Learning (ICL), using zero-shot or few-shot prompts to guide general-purpose LLMs for specific scoring tasks. Relying solely on ICL for complex, project-level evaluation often leads to inaccurate assessments. General LLMs struggle to correctly interpret intricate execution logs or subtle file differences, resulting in significant misalignment with human judgments. For instance, the LLM judger introduced in PaperBench starace2025paperbench achieves a maximum human alignment rate of only 83%. Thus, achieving both evaluation flexibility and high scoring accuracy remains a critical challenge.
To address the two aforementioned challenges, we propose an agent-driven construction approach that enables the creation of project-level benchmarks with flexible metrics at low human cost. Specifically, a state-of-the-art code agent is used to generate both the project scaffolding and a Product Requirement Document (PRD) along with an executable criteria scheme. Human annotators are only required to verify whether the criteria scheme is compatible with the scaffolding interfaces and whether the expected outputs are reasonable, without the need to manually create detailed evaluation standards or reference solutions. This significantly reduces annotation complexity: for PRDBench, annotators with undergraduate-level knowledge in software engineering related fields are able to complete the annotation, with an average of only eight hours needed to finish the scaffolding and metrics for each project, greatly improving annotation efficiency.
Based on this approach, we build a Product Requirement Document (PRD)-centered benchmark, named PRDBench [^1] Source of PRDBench is available in https://github.com/AGI-Eval-Official/PRDBench.. PRDBench consists of 50 real-world coding tasks, each defined by a structured PRD, a well-specified and verifiable criteria scheme, and a standard solution code repository. The coding tasks are sourced from real-world requirements, academic projects, and thesis work, spanning 20 common subdomains. For each task, the PRD provides a criteria scheme that facilitates comprehensive human-like quality assurance (QA) checks. We present a comparison between PRDBench and previous code agent benchmarks in Table 1, where PRDBench offers a comprehensive, multifaceted, project-level benchmark for code agents. During evaluation, we employ the Agent-as-a-Judge paradigm to score code according to the criteria scheme, enabling the evaluation of a wide range of test types beyond unit tests. However, relying on off-the-shelf general LLMs as judges often leads to instability and hallucinations in complex project-level assessments. To ensure highly reliable evaluation, we further introduce a specialized, fine-tuned PRDJudge based on Qwen3-Coder-30B. By training on a curated dataset of evaluation trajectories, our dedicated PRDJudge overcomes the limitations of general models, achieving an over 90% human alignment rate in fixed-interface scenarios.
In summary, our contributions include:
  • We design an agent-driven data production pipeline, where human supervision guides agents to efficiently generate challenging test cases that go beyond current agents capabilities. This pipeline alleviates the need for expertise annotators.
  • We construct PRDBench, which covers 20 common domains in Python software development. Each benchmark item includes a PRD and evaluation metrics, and we design three categories of agent-friendly test types for comprehensive evaluation.
  • We adopt the Agent-as-a-Judge paradigm powered by a specialized fine-tuned model. Based on Qwen3-Coder-30B, our PRDJudge achieves over 90% human alignment in fixed-interface scenarios, enabling flexible, human-like QA assessment beyond traditional unit tests.
  • We conduct extensive experiments on PRDBench, detailing the training process and data construction of the evaluation agent, and providing useful insights into both state-of-the-art code agents and the evaluation paradigm.
Figure 1. Overview of the PRDBench data production workflow. Step1: PRD and Test Plan Initialization; Step 2: Code Scaffold and Criteria Scheme Generation; Step 3: Human Inspection; Step 4: Agent-based Fix and Refinement; Step 5: Remove scaffold.
2. Related Work
2.1. Code Agent Evaluation
To keep pace with the rapid development of LLMs’ and agents’ coding capabilities, recent benchmarks increasingly emphasize the construction of executable and complex tasks within real software projects. These benchmarks are typically built by mining human-generated data from online platforms or through extensive manual annotation. Some benchmarks li2024promptinglargelanguagemodels; huang2024dacodeagentdatascience require annotators to manually create test points for each step, while others jimenez2024swebenchlanguagemodelsresolve; yang2024execrepobenchmultilevelexecutablecode; pan2024codevbenchllmsunderstanddevelopercentric; wu2024repomasterevalevaluatingcodecompletion; hai2025impactscontextsrepositorylevelcode rely on authentic submission records, using the inherent difficulty of submitted code as a basis for evaluation. Additionally, some benchmarks li2024evocodebenchevolvingcodegeneration; zhuo2025bigcodebenchbenchmarkingcodegeneration; peng2024humanevalxlmultilingualcodegeneration leverage existing resources and require expert-driven rewriting, which demands significant manual effort. As code agents become more capable, these annotation paradigms require increasingly specialized expertise, making the recruitment of expert annotators for each domain costly and limiting the scalability and diversity of benchmark construction.
In terms of evaluation metrics, most existing benchmarks adopt relatively narrow criteria, such as unit test pass rates jimenez2023swe; li2024promptinglargelanguagemodels; xiang2025scireplicatebenchbenchmarkingllmsagentdriven, which primarily assess functional correctness. ProjectEval liu2025projectevalbenchmarkprogrammingagents simulates user interaction to evaluate projects, focusing on specific types of tasks. However, these evaluation criteria are limited to specific task types and cannot comprehensively assess diverse, complex software projects.
Figure 2. A task example of PRDBench.
2.2. Agent-as-a-Judge
Due to the inherent limitations of unit test-based benchmarks, recent studies have begun to explore the use of LLMs as judges for agent evaluation, a paradigm often referred to as LLM-as-a-Judge. LLMs can provide capabilities such as image recognition chen2024mllm, code alignment weyssow2024codeultrafeedbackllmasajudgedatasetaligning, and question-answering ho2025llm, breaking the restriction that benchmark metrics must be executable. With the emergence of agents equipped with tools, LLMs are able to read large volumes of files and perform extensive analyses, offering new possibilities for evaluation. DevAI zhuge2024agent is the first work to apply the agent-as-a-judge paradigm to code agent evaluation. It focuses on checking program execution and output format, but lacks comprehensive assessment of overall engineering effectiveness. Some studieswang2025codevisionaryagentbasedframeworkevaluating, have explored LLM-based agent frameworks for evaluating code generation, improving the accuracy of Agent-as-a-Judge through architectural innovations. However, the lack of well-defined metrics and code agents has limited the use of agent-as-a-judge. PRDBench fills this gap with structured criteria and dedicated tools, enabling more robust evaluation.
2.3. Fine-tuned LLMs as Judges
Recent studies suggest that for most tasks, evaluating outputs is inherently easier than generating them together2024finetuning, motivating a line of work that fine-tunes smaller open-source LLMs as dedicated judges. PandaLM wang2024pandalmautomaticevaluationbenchmark, Prometheus kim2024prometheusinducingfinegrainedevaluation, JudgeLM zhu2025judgelmfinetunedlargelanguage, and Auto-J li2023generativejudgeevaluatingalignment demonstrate that fine-tuned models of 7B–33B parameters can match or even surpass GPT-4 on targeted evaluation tasks, while offering significant advantages in cost, speed, and privacy. Notably, Huang et al. huang2025empiricalstudyllmasajudgellm further reveal that such fine-tuned judges essentially operate as highly effective task-specific classifiers, consistently achieving state-of-the-art performance on in-domain benchmarks. These findings motivate our approach of fine-tuning a dedicated PRDJudge specifically for project-level code evaluation, where domain-specific expertise is critical for accurate and reliable scoring.
3. PRDBench
3.1. Seed Tasks
Our seed tasks are carefully curated from three primary sources to ensure diversity and real-world relevance: (1) real-world user development requests collected from an internal AI product development platform; (2) publicly available final projects from Computer Science (CS) courses hosted on GitHub; and (3) code reproduction tasks derived from academic theses in the CS domain. To ensure the suitability and consistency of tasks for PRDBench, we apply a rigorous filtering process. Specifically, we require that each selected task can be fully implemented in Python and that all its associated datasets are publicly accessible. We choose Python as the programming language for PRDBench primarily due to its versatility and comprehensive ecosystem, which supports a wide range of programming paradigms and application domains. This ensures that PRDBench tasks are representative of diverse, real-world scenarios.
3.2. Agent-Driven Data Production
Figure 1 illustrates the agent-driven data production workflow of PRDBench. Throughout the annotation process, we apply SOTA code agents to generate the project scaffold and criteria scheme. Human annotators are only required to supervise the quality of criteria scheme, i.e., whether the criteria aligns with the scaffold interfaces and whether the expected outputs meet the requirements specified in the PRD. It is worth mentioning that manual metric annotation and code modification are not necessary, which greatly reduces both the complexity and time required for annotation.
Step 1: PRD and Test Plan Initialization. After selecting seed tasks, we utilize code agents (such as Claude Code) to generate detailed and standardized PRD documents, ensuring clarity and completeness in task specification. The PRD serves as the evaluation blueprint for PRDBench, and includes sections such as Requirement Overview, Functional Requirements, and Data Requirements. Based on the PRD, we employ GPT-4.1 to generate a corresponding metric outline, structuring test cases using the Arrange-Act-Assert (AAA) methodology wei2025developers. The Arrange step sets up the test case by preparing the necessary files, input data, and environment configurations. The Act step focuses on the core behavior to be tested, such as running the program or performing interaction tests to obtain output results. The Assert step verifies the expected outcomes by checking the system’s response or state, ultimately determining whether the test passes or fails. This structure is applicable to a wide range of code-related testing scenarios, providing PRDBench with a executable testing plan.
Figure 3. Domain distribution of PRDBench.
Step 2: Scaffold and Criteria Generation. After initializing with the PRDs and metric outlines, we employ state-of-the-art code agents to generate the code scaffold for each task. The scaffold includes both module design and interface design, serving as the core framework of the entire project. In some cases, particularly for complex algorithms, the implementation may not be fully correct as long as it does not affect the annotation process. Building upon the scaffold, we further utilize code agents to expand and refine the metric outline into a specific criteria scheme, which includes necessary test interfaces and test artifacts. The presence of well-defined code interfaces in the scaffold simplifies and standardizes the generation of the criteria scheme, making the process more efficient and consistent.
Step 3: Human Inspection. With the scaffold and the criteria scheme in place, the inspection process for human annotators becomes straightforward. Annotators conduct inspection only by running the tests to verify whether the interfaces function correctly and whether the expected outputs in the criteria scheme align with the PRD requirements. Notably, even for tasks sourced from expertise projects or academic papers, annotators only need basic computer science knowledge to perform effective annotation.
Step 4: Agent-Based Fix and Refinement. If any issues are identified during step 3, annotators shall provide targeted feedback to the code agent for refining the scaffold or criteria scheme. The code agent then revises the relevant components, and the inspection process is repeated. This iterative process continues until all issues are resolved. To ensure the rigorous quality and sufficient complexity of our benchmark, we establish a strict inclusion criterion: only tasks that have undergone at least 5 rounds of such iterative human-agent refinement are retained.
Step 5: Remove Scaffold. Finally, the scaffold is removed, retaining only the PRD, criteria scheme, test artifacts, and data. This forces the evaluated agents to generate code from scratch, enabling a rigorous assessment of their end-to-end development capabilities.
An overall example of a PRDBench task is illustrated in Figure 2. Each PRDBench testcase consists of a PRD document and an evaluation suite, which includes a criteria scheme and test artifacts.
3.3. Data Statistics
PRDBench comprises 50 project-level Python tasks that span 20 distinct application domains. To ensure broad and balanced coverage, we utilized GPT-4 to assign domain labels to all candidate tasks before strategically sampling the final dataset. As illustrated in Figure 3, we include at least one representative task from less common domains. For more prevalent areas, such as data processing and machine learning, we select tasks that demand a diverse array of technical approaches and sub-skills. This sampling strategy ensures that PRDBench accurately represents diverse, real-world Python development scenarios.
Figure 4. Scaffold line counts distribution.
These tasks exhibit substantial complexity. On average, each structured PRD contains 105.22 lines of detailed requirements. Furthermore, Figure 4 presents the distribution of scaffold line counts recorded during the annotation process, highlighting the realistic scale of the projects.
For comprehensive evaluation, PRDBench encompasses a total of 1,258 evaluation metrics. These metrics collectively simulate the different stages of quality assurance (QA) in real-world software engineering and are categorized into three main types:
  • Unit Test (408 metrics): As in previous benchmark studies jimenez2023swe, unit tests remain the most effective approach for verifying the functional correctness of individual components and modules. PRDBench retains this rigorous testing paradigm and provides auxiliary pytest-based test scripts.
  • Shell Interaction (732 metrics): For tasks involving command-line interaction or external system operations, predefined shell commands are executed to compare actual outputs against expected results. This ensures the code correctly handles system-level operations and user inputs. PRDBench supplies comprehensive simulated user input files and program entry commands for this purpose.
  • File Comparison (118 metrics): For project-level tasks that generate files or require specific directory structures, the produced artifacts are compared against reference solutions to verify correctness in content, format, and organization. PRDBench provides the raw data for target file generation, the corresponding Python commands, and the reference solution files.
4. PRDJudge
In this section, we present PRDJudge, a specialized evaluation agent developed to automatically and reliably assess code submissions on PRDBench. We elaborate on its tool-augmented evaluation framework, and the fine-tuning methodology designed to align its scoring with human standards.
4.1. Evaluation Framework
Figure 5. Overview of PRDJudge. PRDJudge executes tests based on the criteria scheme using various tools, compares outputs (files or results) with expected outputs, and generates a report for the submitted code.
To automate the validation of code generated by code agents, we introduce PRDJudge, a lightweight evaluation agent tailored for the predefined criteria schemes of PRDBench (illustrated in Figure 5). For each item in the criteria scheme, PRDJudge leverages the executable terminal commands and expected outputs provided by the benchmark to conduct systematic verification.
To interact effectively with the testing environment, PRDJudge is equipped with six core tools, including file reading and writing, command-line execution, image handling, and a specialized judge tool. Among these, the deal graph tool offers a multimodal LLM interface (powered by GPT-4o in this paper), allowing the agent to process visual outputs by submitting both the image and a descriptive prompt to obtain verification results—this is particularly useful for tasks requiring UI or plot validation. Furthermore, the judge tool is specifically designed for PRDBench. It accepts files containing simulated user inputs, enabling the generation of comprehensive terminal logs for thorough analysis without the need for manual input specification.
By leveraging this toolset, PRDJudge automatically adapts to specific code repositories and seamlessly executes the three categories of tests (Unit Test, Shell Interaction, and File Comparison) defined in Section 3.3. Since all supporting files and test categories are self-contained within PRDBench, the complexity of evaluation is greatly reduced. PRDJudge only needs to execute each test, analyze the execution logs or file differences, and generate a detailed report on the performance and correctness of the code agent submissions. However, accurately interpreting complex execution logs and aligning the final scores with human quality assurance standards requires exceptional reasoning capabilities, which motivates the training of a specialized evaluation model, as detailed in the following section.
4.2. Fine-tuning PRDJudge
4.2.1. Human Annotation Protocol
To establish a high-quality ground truth for both training data filtering and final evaluation, we employ a rigorous human annotation process. Detailed information of our annotators is provided in Appendix A. Each scoring point in the benchmark is independently evaluated by two human annotators. In cases of disagreement, a third senior annotator serves as an arbitrator to make the final decision. The annotation follows a strict three-tier ordinal scoring system:
  • Score 2 (Pass): The code executes perfectly, and the output strictly matches the expected requirements specified in the criteria scheme.
  • Score 1 (Partial): The code runs successfully (i.e., the interfaces are correct and no execution crashes occur), but the final output or behavior does not fully align with the expected results.
  • Score 0 (Fail): The code fails to execute entirely, typically due to syntax errors, missing dependencies, or severe runtime exceptions.
4.2.2. Training Data Construction
The training data for PRDJudge consists of high-quality evaluation trajectories. To collect these trajectories, we deploy a highly capable model, Qwen3-Coder-480B-A3B qwen3technicalreport, as the initial judge within the PRDJudge framework. We collect evaluation trajectories from the outputs of 11 different code agents previously tested on PRDBench. To ensure diversity in the training distribution, we randomly sample 8 repositories for each agent to form the training pool.
Crucially, to prevent the model from learning hallucinated or flawed reasoning, we apply a strict two-stage filtering mechanism. Initially, Qwen3-Coder-480B-A3B generated a total of 2,147 evaluation trajectories. In the first stage, we only retain the 1,742 trajectories where the final score predicted by the model exactly matches our human-annotated ground truth. It is worth noting that while the generation of PRDBench tasks is highly efficient and scalable (as discussed in Section 1), establishing this ground truth required a substantial human annotation effort. However, we consider this intensive labeling process a necessary, one-time investment to distill a reliable automated judge, which subsequently enables zero-cost, scalable evaluations for any future code agents. Furthermore, to ensure the efficiency and accuracy of the agent’s testing trajectories, we apply an additional rule-based filter to eliminate trajectories that are excessively long or exhibit invalid tool usage. Although these discarded trajectories achieved the correct final score, their inaccurate tool usage could introduce noise into the evaluation agent’s scoring behavior. Ultimately, this rigorous filtering process yields a final dataset of 911 high-quality training trajectories for fine-tuning.
5. Experiments and Results
5.1. Code Agents
Table 2. Agent Specifications and Open-Source Status.
Agent Type Agent Framework Model Agent OSS LLM OSS
Minimal Basic code tools Qwen3-Coder-480B-A3B (qwen3technicalreport)
GPT-5.2 gpt5.2
Claude-4.5-Sonnet anthropicclaude45
Gemini-3-pro gemini3
Kimi-K2-Instruct-0905 kimiteam2025kimik2openagentic
DeepSeek-V3.2
GLM-4.7 5team2025glm45agenticreasoningcoding
Minimax-M2 minimaxm2
Commercial Gemini CLI mullen2025geminicli Gemini-2.5-pro gemini25
Claude Code anthropic2025claude37 Claude-4.5-Sonnet anthropicclaude45
CodeX codex GPT-5 gpt5
Qwen Code qwencode Qwen3-Coder-480B-A3B qwen3technicalreport
EvalAgent PRDJudge Qwen3-Coder-30B-A3B qwen3technicalreport
We evaluate two categories of code agents in our experiments:
  1. (1)
    Minimal Agents [^2] Code is available in https://github.com/AGI-Eval-Official/Minimal-CodeAgent., implemented using the Agent Development Kit (ADK). These agents are equipped with essential tools for file manipulation, bash scripting, and Python execution. We integrate state-of-the-art LLMs (Qwen3-Coder-480B-A3B, GPT-5.2, Claude-4.5-Sonnet, Gemini-3-Pro, Kimi-K2-Instruct-0905, DeepSeek-v3.2, GLM-4.7, and Minimax-M2) to systematically assess their core capabilities in strategy implementation. Hereafter we denote the models as Qwen3-Coder, GPT-5.2, Claude-4.5, Gemini-3-Pro, Kimi-K2, DeepSeek-v3.2, GLM-4.7, and Minimax-M2 respectively for simplicity.
  2. (2)
    Commercial Code Agents, including advanced CLI-based agents (Claude Code, CodeX, Gemini CLI, and Qwen Code). These agents offer enhanced integration with command-line interfaces and external tools, representing the current state-of-the-art in code agent development.
Each commercial agent and minimal agent utilizes a corresponding backbone LLM. Commercial agents typically benefit from vendor-specific fine-tuning and optimization, resulting in more stable and robust performance within their respective frameworks. While minimal agents are designed to facilitate fair comparison across different LLMs, commercial agents provide an assessment of the latest advancements in agent capabilities. The mapping between agent frameworks and backbone models is summarized in Table 2.
5.2. Experimental Setups
For code agents, we select the latest model from each family that is compatible with the agent framework. For all LLMs used in our work, we set the temperature to be $0.1$, and the max tokens identical to their official APIs’ setting. We set top-p to $1.0$, top-k to $100$, and presence penalty to be default to the API. Each code agent is executed using a Python virtual environment that contains necessary and useful packages for the agents. If the code agent need other packages, we allow it use pip to install them.
For PRDJudge training, we adopt Qwen3-Coder-30B qwen3technicalreport as our base model, balancing strong coding capabilities with deployment efficiency. To optimize training costs while maintaining performance, we employ Low-Rank Adaptation (LoRA) for parameter-efficient fine-tuning. The LoRA rank is set to 16. The model is trained for 1 epoch using a learning rate of $2.0\times 10^{-4}$ and a warmup ratio of 0.03. The fine-tuning process is conducted on a cluster of 8 NVIDIA H800 GPUs.
5.3. PRDJudge Evaluation
5.3.1. Evaluation Setup
To evaluate the reliability and generalizability of PRDJudge, we construct two complementary test sets with full human annotation. (1) In-Domain Test Set. We collect code submissions generated by the same 11 code agents used in the training data construction phase. For each agent, we randomly sample 2 repositories (different to training data), yielding a total of 513 scoring metrics for human annotation. This test set is designed to assess PRDJudge’s alignment with human judgment under familiar agent and repository distributions. (2) Out-of-Domain Test Set. To evaluate the generalization capability of PRDJudge beyond the training distribution, we additionally construct an out-of-domain test set, where the code agents whose outputs are evaluated did not appear in the training data. This set comprises 26 repositories spanning diverse domains, resulting in a total of 605 scoring metrics. This setting more closely reflects real-world deployment scenarios, where PRDJudge may encounter previously unseen agent behaviors and implementation styles.
All annotations undergo a rigorous two-round labeling process to ensure quality and consistency. Specifically, each metric is independently annotated by two annotators. For cases where both annotators reach the same score, we randomly sample 10% of such instances for a third-round quality check by a dedicated inspector. For cases where the two annotators disagree, the inspector reviews both annotations and provides a final adjudicated score, taking into account the reasoning of both parties. This multi-stage annotation pipeline ensures high inter-annotator reliability and minimizes subjective bias in the ground-truth labels.
5.3.2. Evaluation Metric
To assess the reliability of PRDJudge against human standards, we employ the Human Alignment Rate (HAR) as our primary evaluation metric. HAR calculates the exact match accuracy between the scores predicted by the evaluation model and the human-annotated ground truth. It intuitively reflects the overall correctness and reliability of the judge.
Table 3. Performance comparison of PRDJudge and baseline models on the In-Domain and Out-of-Domain test sets. The accuracy is measured by the Human Alignment Rate (HAR, %). Avg. Token and Avg. Time (in seconds) denote the average context length and inference time required per metric evaluation.
Model In-Domain Out-of-Domain Avg. Token Avg. Time
Human 95.83 95.83 - ¿ 600
Qwen3-Coder (480B) 90.32 87.91 91,985 114.96
Claude-4.5 90.64 88.10 166,386 174.64
GPT-5.2 89.76 87.09 33,182 133.96
Qwen3-Coder (30B) 70.52 55.46 151,411 162.34
PRDJudge (ours) 91.75 92.69 103,712 107.85
5.3.3. Human Alignment
Table 3 presents the evaluation results against human annotations, leading to three key conclusions: (1) PRDJudge achieves state-of-the-art human alignment. PRDJudge outperforms all baseline models, including massive proprietary LLMs, and approaches the human inter-annotator agreement upper bound. The massive improvement over its base model (Qwen3-Coder 30B) highlights the necessity of specialized training for complex, long-context PRD-based evaluation. (2) PRDJudge exhibits exceptional out-of-domain generalizability. Zero-shot baseline models experience a noticeable performance drop on the out-of-domain test set, indicating this set is intrinsically harder (likely due to complex edge cases from unseen, stronger agents). Despite this challenge, PRDJudge maintains highly stable performance. This proves it genuinely internalizes the underlying evaluation criteria rather than merely overfitting to the training agents’ specific error patterns. (3) PRDJudge offers superior computational efficiency. Alongside the highest accuracy, PRDJudge records the lowest average inference time. Compared to its base model, it significantly reduces token consumption and inference time, suggesting that specialized training enables more concise, targeted reasoning. This makes PRDJudge a highly practical and scalable solution for real-world automated evaluation.
5.3.4. Stability Analysis
To evaluate scoring stability, we conduct three independent evaluation runs for each metric and measure the Unanimous Agreement Rate (UAR) and Pairwise Agreement Rate (PAR). Detailed results are provided in Appendix B. PRDJudge demonstrates the highest stability among all evaluated models, achieving a UAR of 94.19% and a PAR of 96.07%. The superior consistency indicates that our specialized fine-tuning helps PRDJudge internalize a deterministic evaluation mapping from PRDs to code, significantly mitigating the reasoning variance and generation randomness typically observed in general-purpose LLMs.
5.3.5. Case Study
To intuitively understand why PRDJudge achieves superior performance and efficiency, we trace and analyze the tool-use trajectories of different models during the evaluation process. The complete tool usage distribution and case study is provided in Appendix C. First, PRDJudge follows a logical explore-then-evaluate workflow, prioritizing repository exploration and file reading to understand code logic before making judgments. In contrast, GPT-5.2 exhibits a blind action pattern, characterized by excessive command execution and file writing with minimal code inspection. Second, PRDJudge demonstrates high precision. While its base model (Qwen3-30B) suffers from tool spamming due to convoluted reasoning, PRDJudge requires significantly fewer, more targeted calls than even Claude-4.5. This confirms that specialized training enables the model to optimally gather evidence rather than merely increasing tool-use frequency.
5.4. Code Agent Evaluation
5.4.1. Overall Results
Table 4 presents agent performances across two phases: Round 1 (DEV) assesses zero-shot project implementation based on PRDs, while Round 2 (DEBUG) evaluates iterative refinement using PRDJudge feedback. Our main findings are: (1) Foundation models dictate zero-shot development ceilings. Frontier models (e.g., Claude-4.5, GPT-5.2) consistently dominate the DEV phase, indicating a strong correlation between intrinsic LLM reasoning capacity and the ability to generate project-level scaffolds from scratch. (2) Commercial frameworks constrain initial generation but excel in iterative debugging. Minimal agents generally outperform their Commercial counterparts in DEV, suggesting that complex tool-use guardrails in commercial frameworks restrict zero-shot generation. Conversely, commercial agents demonstrate superior debugging capabilities, effectively utilizing feedback for targeted fixes. (3) Unconstrained models risk performance degradation during long-context debugging. High-performing Minimal agents (e.g., Claude-4.5, DeepSeek-V3.2) experience substantial drops in the DEBUG phase. Without robust agentic scaffolding to manage global states, attempting localized modifications frequently breaks previously functioning modules, underscoring the necessity of structured designs for safe code maintenance. (4) LLMs exhibit divergent generation and refinement profiles. Within Minimal agents, we observe three archetypes: (i) strong zero-shot generation but high variance in context maintenance, causing debugging regressions (Claude-4.5, DeepSeek-V3.2); (ii) limited initial generation but strong instruction-following for localized fixes, yielding significant improvements (Kimi-K2, Gemini-3-Pro); and (iii) consistent performance across both phases (GPT-5.2, Qwen3-Coder). This suggests future agent designs must adapt to the specific reasoning characteristics of their backbone models.
Table 4. Average pass rate of code agents on PRDBench (in %). The best results are highlighted in bold, and the second-best results are underlined.
[…]
Minimal GPT-5.2 62.49 69.00 +6.51
Claude-4.5 69.19 56.40 -12.79
Gemini-3-Pro 22.76 27.28 +4.52
Qwen3-Coder 43.84 48.29 +4.45
Kimi-K2 20.52 36.17 +15.65
DeepSeek-V3.2 40.11 24.80 -15.31
GLM-4.7 38.39 35.33 -3.06
Minimax-M2 17.60 24.75 +7.15
Commercial CodeX 62.09 65.02 +2.97
Claude Code 56.65 70.25 +13.60
Gemini CLI 11.29 21.51 +10.22
Qwen Code 39.91 35.95 -3.96
Figure 6. Error rates of code agents on different types of test cases.
5.4.2. Cost of Code Agents
Appendix D details the computational expenditures of the evaluated LLM-based code agents across two experimental iterations, revealing distinct operational patterns. First, the generation and maintenance profiles diverge significantly across agent configurations. The execution of minimal LLM agents is characterized by high-volume initial code generation, followed by high volatility during the debugging phase, where modifications vary from minor syntax corrections to extensive logic restructuring within the context window. Conversely, commercial LLM frameworks operate under highly constrained generation parameters. Their maintenance phase is defined by localized modifications, which restricts code alterations to a stable and minimal scope regardless of task complexity. Second, the underlying drivers of computational overhead differ fundamentally. For minimal agents, token consumption and inference latency scale directly with the volume of generated code. In contrast, commercial agents incur substantial token expenditures and latency that remain largely decoupled from the actual volume of code modification. This resource distribution reflects an architectural reliance on complex tool-use reasoning, continuous environment synchronization, and iterative state management rather than baseline zero-shot drafting speed.
5.4.3. Error Analysis of Code Agent
Figure 6 shows the error rates of code agents across different types of test cases. The error rate reflects the proportion of failed test cases within each category for the code generated by the agents. This comparison highlights the strengths and weaknesses of various agents when handling diverse evaluation scenarios. We observe two main findings from the results. First, there is no single category of test cases that is consistently easy for code agents. The error rates are relatively uniform across the three types of test cases, indicating that PRDBench’s test case design is well-balanced and effectively covers a broad spectrum of code implementation scenarios. Second, unit test cases are noticeably more challenging to debug compared to the other two types. Most code agents fail to resolve these issues even after the debugging phase. This is primarily because unit test cases require agents to read and understand the test function code before making corrections, whereas the other two test types only require agents to compare actual and expected outputs, which lowers the reasoning complexity.
5.4.4. Free Development
Figure 7. Comparison of code agent scores under fixed inference and free development modes on PRDBench.
In real-world scenarios, developers often build from scratch based solely on a PRD. Thanks to its scaffold-removal design, PRDBench inherently supports this unconstrained free development setting. To explore this, we prompted PRDJudge to dynamically adapt test interfaces to the agents’ free-form code.
Figure 7 compares agent performance across both modes. While relative rankings remain stable, the performance variance decreases significantly under free development. This highlights a key confounding variable: it is difficult to disentangle agent logic errors from PRDJudge’s failures in dynamically rewriting test interfaces. Due to this evaluation noise, we conclude that Fixed Inference remains the most reliable benchmarking setting at present. The free development mode serves as a proof-of-concept, underscoring the inherent difficulties of unconstrained evaluation that we aim to address in future work.
6. Conclusion
In this paper, we introduce PRDBench, a comprehensive benchmark for evaluating repository-level code agents against Product Requirement Documents (PRDs). To address the scalability limitations of human evaluation and the inherent inefficiencies of general-purpose LLM judges, we propose PRDJudge, a specialized evaluator fine-tuned for code-to-PRD alignment. PRDJudge achieves state-of-the-art human alignment and exceptional out-of-domain generalizability. Trajectory analysis demonstrates that PRDJudge adopts a systematic verification strategy, effectively mitigating the redundant tool invocations and execution-heavy inefficiencies prevalent in general LLMs, thereby providing a robust evaluation infrastructure. Leveraging this framework, our extensive evaluations on PRDBench reveal critical insights into current code agents. We observe that while the intrinsic capabilities of foundation models determine the ceiling for zero-shot development, commercial frameworks constrain initial generation but significantly enhance iterative debugging. Furthermore, unconstrained LLMs frequently introduce regressions during long-context refinement, and system-level interactions pose far greater challenges than isolated unit tests. Ultimately, by shifting the focus from isolated functional correctness to comprehensive PRD alignment, this work establishes a scalable and robust evaluation paradigm, providing a critical foundation for the future development and assessment of code agents.
Future Work. While PRDBench excels in fixed-interface settings, evaluating unconstrained Free Development introduces noise when dynamically adapting test scripts to unpredictable code. Future work will focus on enhancing PRDJudge’s dynamic interface-mapping capabilities to achieve reliable automated evaluation in fully unconstrained scenarios.
{acks}
The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62322603, 62177033, 62502310). We extend our sincere gratitude to Shuang Zhou, Yuan Zhang (张元), Yuan Zhang (张源), Ying Xie, Hao Zheng, Yan Wang, Zhiqiang Li, Cheng Li, Bangpeng Wei and Huaying Xia for their contributions to data annotation.
Appendix A Recruitment and Compensation of Annotators
We employ a total of 8 annotators, including five full-time annotators and three part-time annotators. We list their educational background in Table 5.
Among them, F1 to F6 and P1 participants in PRDBench construction and each of them contribute at least 5 problems.
F1, F2, P2 and P3 participants in data examination.
F1 to F6, P2 and P3 participants in human score annotation to evaluate and train PRDJudge.
Table 5. Background of annotators.
Type ID Major Degree Years of Experience as Developer
Full-time F1 Computer Science Master 10
Full-time F2 Software Engineering Master 6
Full-time F3 Applied Statistics Master 1
Full-time F4 Health Service and Management Bachelor 2
Full-time F5 Information Management Bachelor 4
Full-time F6 Economics and Trade Management Bachelor 6
Part-time P1 Computer Science Bachelor 9
Part-time P2 Computer Science Senior Undergraduate -
Part-time P3 Computer Science Senior Undergraduate -
All annotators receive daily wages (part-time) or monthly salaries (full-time) according to local labor regulations.
Appendix B Stability Analysis of PRDJudge
In this section, we provide the detailed evaluation metrics and the complete results for the stability analysis discussed in the main text.
B.1. Metrics Definition
Due to the inherent probabilistic nature of Large Language Models, their generated evaluations can exhibit variance across different runs, especially when dealing with long-context reasoning tasks like PRD-based code evaluation. To rigorously quantify scoring stability, we conduct three independent evaluation runs for each metric and employ two criteria:
  • Unanimous Agreement Rate (UAR): This metric calculates the percentage of evaluation instances where the model predicts the exact same score across all three independent runs. Let $s_{1},s_{2},s_{3}$ be the predicted scores in three runs. The UAR for a single metric is 1 if $s_{1}=s_{2}=s_{3}$, and 0 otherwise. It represents the strictest measure of consistency.
  • Pairwise Agreement Rate (PAR): This metric computes the average agreement rate across all three possible pairs of the runs (i.e., Run 1 vs. Run 2, Run 2 vs. Run 3, and Run 1 vs. Run 3). It provides a more granular view of stability, accounting for partial agreements even when unanimous agreement is not reached.
B.2. Results and Discussion
As shown in Table 6, evaluating code based on complex PRDs is a highly challenging task that can induce reasoning fluctuations in general-purpose LLMs. Even state-of-the-art proprietary models like Claude-4.5 and GPT-5.2 exhibit noticeable variance, failing to reach unanimous agreement in approximately 7% to 9% of the cases. This variance often stems from the models attending to different constraints within the lengthy PRD context across different runs. In contrast, PRDJudge achieves the highest stability with a UAR of 94.19% and a PAR of 96.07%. Through specialized fine-tuning on domain-specific evaluation trajectories, PRDJudge has learned a more robust and deterministic mapping between PRD requirements and code implementations. This significantly mitigates the generation randomness, making PRDJudge a more reliable evaluator for automated pipelines where consistent feedback is critical.
Table 6. Stability analysis of different models across three independent runs. UAR denotes Unanimous Agreement Rate, and PAR denotes Pairwise Agreement Rate. Higher values indicate better scoring stability.
Model UAR (%) $\uparrow$ PAR (%) $\uparrow$
Qwen3-Coder (480B) 91.95 94.57
Claude-4.5 93.07 95.32
GPT-5.2 91.22 94.14
Qwen3-Coder (30B) 70.92 79.42
PRDJudge (ours) 94.19 96.07
Appendix C EvalAgent Analysis
C.1. Tool Use Distribution
Figure 8 illustrates the distribution and total frequency of tool calls made by different models during the evaluation process. This behavioral analysis provides deeper insights into the reasoning strategies adopted by each evaluator.
Figure 8. Tool Use Distribution of Agents.
The heatmap highlights several critical anomalies in general-purpose LLMs when confronted with complex, repository-level evaluation tasks:
  • GPT-5.2’s Skewed Trajectory: GPT-5.2 exhibits a highly irregular pattern. It rarely uses exploration tools like list workspace (1.85%) and read file (4.85%). Instead, it heavily relies on write file (40.80%) and run system command (28.64%). This suggests that GPT-5.2 attempts to solve the evaluation by aggressively executing trial-and-error operations rather than systematically reading the codebase.
  • Qwen3-Coder (30B)’s Inefficiency: The base model, Qwen3-30B, records an exorbitant 7,035 total tool calls. This massive volume indicates that the model frequently falls into reasoning loops or hallucinates incorrect tool parameters, forcing it to retry repeatedly without progressing toward a final judgment.
Conversely, PRDJudge demonstrates an optimal balance. It dedicates the highest proportion among all models to repository exploration (list workspace, 20.26%), ensuring a solid grasp of the project structure. Its total call volume (2,972) is less than half of its base model’s, proving that the fine-tuning process effectively instills a disciplined, evidence-driven evaluation methodology.
C.2. Case Study
In this section, we provide detailed tool-use trajectories to illustrate the behavioral differences between PRDJudge and baseline models. These cases highlight the logical exploration, decision precision, and control stability of our model.
C.2.1. Case Study I: Logical Exploration (Explore-then-Evaluate)
Metric: 0.1.1 Environment and Documentation: Provide clear documentation (README.md)
Repository: ADK_claude_3_7_1
PRDJudge: 2
As shown in the trajectory below, PRDJudge demonstrates a systematic approach to documentation verification. Unlike baselines that failed to locate the relevant files, PRDJudge first maps the workspace and then cross-references the actual file with the benchmark’s expected template.
PRDJudge Trajectory (Case I) (1) list workspace: Enumerate root directory to identify src/ and evaluation/ folders. (2) read_file(src/README.md): Inspect the project’s documentation for setup and run commands. (3) read_file(evaluation/expected_README.md): Retrieve the evaluation checklist/template. (4) write file: Output a JSON report confirming the README satisfies all PRD requirements. (5) exit loop: Terminate after successful evidence collection.
Behavioral Contrast: PRDJudge reaches the correct judgment purely via static reasoning over documentation. It avoids unnecessary system commands, showing a tight evidence chain between the PRD requirement and the source artifacts.
C.3. Case Study II: Blind Action vs. Grounded Review
Metric: 2.1.4b CSV Batch Import
Repository: ADK_gpt_5_4
Scores: Human: 2 — GPT-5.2: 1 — PRDJudge: 2
Analysis.
This case highlights the Blind Action failure mode in general LLMs. GPT-5.2 acts as a tester rather than a reviewer, relying solely on execution logs without understanding the underlying code or specification.
Table 7. Comparison of Tool-use Trajectories for Case II.
Step GPT-5.2 (Blind Action) PRDJudge (Grounded Review)
1 judge: Run main.py with input script. list workspace: Map project structure.
2 write file: Assign Score 1 (log mismatch). read file: Inspect CSV template and PRD.
3 exit loop judge: Verify CLI flow against PRD.
4 - write file: Assign Score 2 (semantic match).
5 - exit loop
Behavioral Contrast: GPT-5.2 suffers from brittleness by checking for literal string matches in terminal logs. Conversely, PRDJudge integrates execution into a broader evidence chain, ensuring the decision is based on semantic alignment with the PRD.
C.4. Case Study III: Tool Spamming and Control Stability
Metric: 0.1.1 Environment and Documentation: Provide a clear
README.md
Repository: ADK_gpt_5_4
Scores: All models agree on Score 2, but efficiency varies significantly.
Analysis.
While both models reach the correct score, the base model (Qwen3-Coder 30B) exhibits severe tool spamming, failing to terminate efficiently in a complex tool environment.
  • Qwen3-Coder (30B) Trajectory (347 messages):
    $\texttt{list workspace}\rightarrow\texttt{read file}\rightarrow\texttt{write file}$
    $\rightarrow\underbrace{\texttt{exit loop}\rightarrow\dots\rightarrow\texttt{exit loop}}_{\text{Repeated 170 times}}$
  • PRDJudge Trajectory (9 messages):
    $\texttt{list workspace}\rightarrow\texttt{read file}\times 2\rightarrow\texttt{write file}\rightarrow\texttt{exit loop}$
Behavioral Contrast: The base model lacks decision determinism, getting stuck in redundant loops. PRDJudge, through specialized fine-tuning, exhibits stable, low-entropy control, collecting exactly the evidence needed and exiting immediately.
Appendix D Cost of Code Agents
Table 8. Cost statistics of code agents. Code indicates the average number of lines generated in Round 1 (development) and the average number of lines modified in Round 2 (debug). Avg. Input and Avg. Output is the average number of token costs of code agents. Relevant log data for CodeX is unavailable, so its results are not reported.
Round 1 (DEVELOPMENT) Round 2 (DEBUG)
Time (s) Avg. Input Avg. Output Code (lines) Time (s) Avg. Input Avg. Output $\Delta$ Code (lines)
Minimal Agent
GPT-5.2 898.05 1614722.40 20243.60 1748.24 906.14 328124.94 5880.08 87.78
Claude-4.5 1986.97 4167292.81 45362.87 3035.36 1387.24 3031463.60 26578.81 1247.45
Gemini-3-Pro 1838.73 2487502.55 57020.49 3149.24 2077.41 2834587.39 62862.83 178.83
Qwen3-Coder 677.83 1973132.60 21926.21 2136.44 544.11 1447580.76 13683.88 217.46
Kimi-K2 1507.05 1241962.54 33457.28 2851.24 2382.83 2040627.56 28273.62 1254.37
DeepSeek-V3.2 3148.31 3114625.31 65983.08 6232.84 1400.76 3733697.44 33686.88 934.06
GLM-4.7 2011.51 3176073.60 41262.10 4132.25 1386.35 3105477.64 31583.20 500.20
Minimax-M2 1238.56 3869757.84 61623.16 2996.82 1411.54 3739755.16 42433.44 276.26
Commercial Agent
CodeX - - - 1907.72 - - - 109.58
Claude Code 2973.0 6469458.40 31599.4 2717.50 1257.51 6695643.42 19727.84 207.38
Gemini CLI 2740.70 834125.00 20238.96 1729.04 2067.61 5776370.08 24611.50 277.80
Qwen Coder 1183.42 3108353.62 31400.28 2610.16 9582.81 988041.65 5604.89 112.41
Table 8 presents detailed cost statistics for each code agent across two phases: Round 1 (Development) and Round 2 (Debug). An analysis of the empirical data yields three primary observations.
  • Divergent Generation Strategies: Minimal agents generally demonstrate a higher volume of initial code generation. For instance, DeepSeek-V3.2 generated an average of 6,232.84 lines in Round 1, significantly exceeding the output of commercial agents. This indicates that minimal agents are often optimized for comprehensive, end-to-end generation in a single pass, whereas commercial frameworks typically employ more modular or constrained protocols.
  • Logic Stability in Debugging: Debugging strategies reveal distinct patterns between the two categories. Minimal agents show significant variability, ranging from localized interventions to extensive architectural restructuring. Conversely, commercial agents adhere to a uniform, conservative protocol characterized by a consistently low volume of code changes. This suggests that commercial frameworks prioritize incremental adjustments to maintain codebase integrity, while minimal agents remain highly dependent on the model’s initial output quality.
  • Decoupling of Time and Modification Scale: For commercial agents, time expenditure operates independently of the scale of modification. While minimal agents’ time costs scale proportionally with output, commercial systems often incur disproportionately high latency during debugging relative to their minimal $\Delta$ Code. For example, the commercial version of Qwen Coder requires 9,582.81 seconds to modify only 112.41 lines, suggesting that temporal costs stem from intensive backend processes like multi-step verification and environment synchronization.
Appendix E Full Prompts
E.1. Prompt of code agents
Prompt for Round 1 (development):
Please develop a complete Python project (ID:{ID}) according to the requirements specified in the project documentation (src/PRD.md), and with reference to the expected test metrics (evaluation/detailed_test_plan.json).
### Requirements
1. Strictly implement all functional requirements described in PRD.md, ensuring that every feature is fully realized and that no requirements are omitted.
2. Closely follow the testing schemes defined in detailed_test_plan.json, ensuring that your implementation process and interfaces fully comply with the testing specifications, so that QA testing can be carried out directly using detailed_test_plan.
3. Submit all project code and related files completely under the src/ directory, ensuring that the project structure is clear and maintainable.
4. Do not ask any intermediate questions during the development process. Complete the entire project and submit directly.
Prompt for Round 2 (debug):
You are provided with the following resources:
- The project requirements and code in src/ (including PRD. md and source code)
- The evaluation criteria and related files in evaluation/
- The development score report of code in src/ is in reports/
Please analyze the items in the score report where points were deducted, and modify the code to address these issues. Ensure that your revised code fully complies with both the evaluation criteria and the requirements specified in PRD.md.
Prompt for free development in Section 4.7:
Please develop a complete Python project (ID:{ID}) according to the requirements specified in the project documentation (src/PRD.md).
### Requirements
1. Strictly implement all functional requirements described in PRD.md, ensuring that every feature is fully realized and that no requirements are omitted.
2. Independently design the implementation details, including process flows and interfaces, based on the requirements in PRD.md, so that each function can be clearly and completely used via the CLI.
3. Submit all project code and related files completely under the src/ directory, ensuring that the project structure is clear and maintainable.
4. Do not ask any intermediate questions during the development process. Complete the entire project and submit directly.
E.2. Prompt of PRDJudge
Prompt for PRDJudge:
# Role & Objective
Evaluate the project implementation at ‘{project_dir}‘ against the metric: ‘{metric_name}‘.
**Core Rule**: Strict adherence to the evaluation criteria. Do NOT modify any project or evaluation files.
# Context & Paths
- **Project Root**: ‘{project_dir}‘ (Source code in src/‘)
- **Evaluation Files**: ‘{project_dir}/evaluation
- **Output Report**: ‘{metric_report_file}‘
- **Working Directory**: {os.path.dirname(project_dir)}. You can use parameter workspace_dir in tools to change cwd of your code execution (Usually is the project directory).
# Evaluation Metric Data
{json.dumps(metric_data, ensure_ascii=False, indent=2)}
- metric’: the metric name
- description’(Important): Arrange-Act-Assert description of the test metric
- type’: the type of the test metric, can be unit_test’, shell_interaction and file_comparison
- testcases (Important): reference execution commands and input files
- expected_output / expected_output_files (Important): expected output after executing the testcases
- input_files’: input files for the testcases, check the file existence in the project directory.
# Standard Execution Trajectory (Strict Sequence)
For the metric above, you **MUST** follow this exact trajectory. Do NOT improvise.
**Step 1: Extraction & Freeze**
- Extract the test_command and test_input from the testcases field.
- **IMMUTABLE RULE**: You are **FORBIDDEN** from modifying the test_command string.
- Do NOT change python to python3‘.
- Do NOT fix apparent typos.
- Do NOT alter relative paths.
- **Execute the command string VERBATIM (exactly as provided).**
**Step 2: Execution (Select Tool)**
- If type is **shell_interaction**:
- Call judge(entry_command=test_command, input_file=test_input, workspace_dir=...)‘.
- If type is **unit_test** or **file_comparison**:
- Call run_system_command(command=test_command, workspace_dir=...)‘.
- *Note*: For file_comparison‘, delete old output files before running.
**Step 3: Verification & Scoring**
- Apply the **Scoring Criteria** below based on the execution result.
# Execution & Scoring Criteria (Strict)
Assess the code based on the type field in the metric data above.
DO NOT modify testcases commands in the metric data. Execute the testcases commands as they are.
**[Global Fatal Errors] (Score 0 immediately)**
1. **Environment/Syntax**: ImportError‘, No module named‘, SyntaxError‘, IndentationError‘.
2. **Runtime Blocks**: TypeError‘, AttributeError‘, NameError (if they prevent the main logic from running).
3. **Status**: "no test run" or Timeout (> 1 min).
4. **Multi-Req**: If a test point has multiple requirements, ONE failure = Score 0.
## [Scoring by Type]
**1. unit_test**
**[Execution Strategy]**
1. Run the provided testcases command (usually pytest‘).
2. Check the terminal output for "FAILURES" or "ERRORS".
3. **Do NOT** manually check code logic; rely on the pytest result.
**[Scoring Rules]**
- **0**: Code broken by [Global Fatal Errors] or meets **interface mismatch** (Function signature mismatch (wrong name/args) causing TypeError/AttributeError).
- **1**: Code runs but fails AssertionError‘, no matter what the reason is.
- **2**: All tests run successfully.
**2. file_comparison**
**[Execution Strategy]**
1. **Clean**: Delete the output file (not expected_output_files) if it already exists (to ensure new generation).
2. **Run**: Execute the testcases command. If Program crashes *before* generating the file, stop the execution and give the score of 0.
3. **Locate**: If the path is not printed, **READ SOURCE CODE** to find where the file is saved.
4. **Compare**: Read the generated file and compare content with expected_output‘.
**[Scoring Rules]**
- **0**: Program crashes *before* generating the file, or file not found.
- **1**: File is generated successfully, but content differs from expected_output‘.
- **2**: File content matches exactly.
**Note1**: If the file is generated and correct, ignore terminal errors occurring *after* generation.
**Note2**: If the file path is not explicitly printed to the terminal, you **MUST** read the sourcecode to locate the generated file and verify it.
**3. shell_interaction**
**[Execution Strategy]**
1. **Run**: Execute the testcases command and use given input to test the program.
2. **Verify Pre-steps**: Check if the program accepts the input without crashing immediately.
3. **Verify Output**: Compare the terminal output (stdout) with expected_output‘.
- **0**:
- Input mismatch (Cannot handle ‘. in file).
- Program crashes *before* producing the target output.
- Pre-steps fail.
- **1**:
- Pre-steps correct (handles ‘. in‘), but final output is wrong/empty.
- **2**:
- Main function output matches expected_output‘.
- **Crucial Exception**:
Check if the content of input_file exists in the judge log, if not, score = 0.(more important than the next rule)
If the expected output is printed, ignore errors occurring *afterwards* (e.g., crash on exit menu), unless description strictly demands a clean exit.
**[Conflict Rule]**
If description conflicts with expected_output‘, **‘expected_output prevails**.
**[Reference Scoring]**
- If the code exits before inputting the test_input, the score is 0.
# Execution Guidelines
1. **Environment**: If you encounter No module named xxx’, try: source activate /mnt/dolphinfs/hdd_pool/docker/user/hadoop-aipnlp/EVA/zhangbolun06/env/evalADK‘.
2. **Tools**:
- Use deal_graph if image analysis is required.
- Use write_file to save the final report.
3. **Working Directory**: You can explicitly pass the workspace_dir parameter to tools to change the execution CWD. (Default is ‘{os.path.dirname(project_dir)}‘).
# Tips
If the code is unable to run, please give the score of 0 and report it in the report.
If the detailed_test_plan mentions that image analysis is required, use the "deal_graph" tool to analyze the images.
Use the write_file tool to write the report content into a file, passing it to the content variable as a string type when writing.
If the metric has more than one testcase, you can choose to use "start_interative_shell" tool to start a new shell session for each testcase. And if you need to input content to the new shell, use the "run_interactive_shell" tool.
# Output Requirement
Save the result to ‘{metric_report_file}‘ in JSON format.
**JSON Structure:**
"metric": "{metric_name}",
"description": "{description}",
"score": <0, 1, or 2>,
"explanation": "Concise reasoning. Mention specific error types (e.g., ImportError) or logic gaps."
# Examples:
<example_1>
<metric_data>
"metric": "Model Visualization Output",
[…]
"type": "file_comparison",
"testcases": [
"test_command": "python -c \"import sys; sys.path.append(’src’); from models.draw import draw_visualization; draw_visualization()\"",
"test_input": null
"input_files": null,
"expected_output_files": [’evaluation/expected_fx.png’],
"expected_output": "The file output/images/fx.png has been successfully created, The image clearly displays f(x) (increases monotonically from 0 to 1)."
</metric_data>
<expected_trajectory>
Model: OK, I will evaluate the project against the metric: Model Visualization Output.
I will execute the testcases command to evalaute the metric. The command is: python -c \"import sys; sys.path.append(’src’); from models.draw import draw_visualization; draw_visualization()\".
[tool_call: list_workspace_files(workspace_dir=’/path/to/project’)].
[tool_call: read_file(file_path="/path/to/project/src/models/draw.py")].
The visulization figure will be saved in the output/images/fx.png file. Now, I will delete the existing file output/images/fx.png first.
[tool_call: run_system_command(command="rm -f output/images/fx.png", workspace_dir=’/path/to/project’)].
Now, I will execute the testcases command and start my evaluation.
[tool_call: run_system_command(command="python -c \"import sys; sys.path.append(’src’); from models.draw import draw_visualization; draw_visualization()\"", workspace_dir=’/path/to/project’)].
The program runs successfully! Then I will check the generated file output/images/fx.png.
Before check the file, I will check the file existence in the project directory.
[tool_call: run_system_command(command="ls -l output/images/fx.png", workspace_dir=’/path/to/project’)].
The file exists. Now, I will examine the generated file.
[tool_call: deal_graph(graph_path_list=["output/images/fx.png"], prompt’:’Please check if the image clearly displays f(x) (increases monotonically from 0 to 1).’)].
deal_graph tool returns the result: The image is clearly displaying f(x), but the f(x) curve is not increasing monotonically. According to the scoring criteria, the file is generated successfully, but content differs from expected_output‘. Thus the score is 1. Now, I will write the report content into a file.
[…]
[tool_call: exit_loop()]
</expected_trajectory>
</example_1>
<example_2>
<metric_data>
"metric": "2.1.2b SIR Model R0 Calculation",
[…]
"type": "shell_interaction",
"testcases": [
"test_command": "python -c \"import sys; sys.path.append(’src’); from models.sir_model import SIRModel; model = SIRModel(); model.run_simulation()\" | grep -E Basic regeneration number|R0’",
"test_input": null
"input_files": null,
"expected_output_files": null,
"expected_output": "During the SIR model operation, the R0 calculation result should be displayed, including the calculation formula R0 = beta/gamma and specific numerical results. If default parameters are used, R0 should show a reasonable value"
</metric_data>
<expected_trajectory>
Model: OK, I will evaluate the project against the metric: 2.1.2b SIR Model R0 Calculation. This is a shell interaction metric.
I will execute the testcases command to evalaute the metric.
[tool_call: run_system_command(command="python -c \"import sys; sys.path.append(’src’); from models.sir_model import SIRModel; model = SIRModel(); model.run_simulation()\" | grep -E Basic regeneration number|R0’", workspace_dir=’/path/to/project’)].
The program runs successfully! Then I will check the terminal output. The terminal outputs an error No module named SIRModel’. According to the scoring criteria, the program crashes *before* producing the target output. Thus the score is 0. Now, I will write the report content into a file.
[…]
[tool_call: exit_loop()]
</expected_trajectory>
</example_2>
<example_3>
<metric_data>
"metric": "Pytest Example",
[…]
"type": "unit_test",
"testcases": [
"test_command": "cd evaluation && python -m pytest tests/pytest_example.py::TestPytestExample::test_table_display_data_integrity_validation -v",
"test_input": null
"input_files": null,
"expected_output_files": null,
"expected_output": "The pytest shall display the test status as PASSED’."
</metric_data>
<expected_trajectory>
Model: OK, I will evaluate the project against the metric: Pytest Example. This is a unit test metric.
I will execute the testcases command to evalaute the metric.
[tool_call: run_system_command(command="cd evaluation && python -m pytest tests/pytest_example.py::TestPytestExample::test_table_display_data_integrity_validation -v", workspace_dir=’/path/to/project’)].
The program runs successfully, but the test fails. According to the scoring criteria, the program runs successfully, but the test fails. Thus the score is 1. Now, I will write the report content into a file.
[…]
[tool_call: exit_loop()]
</expected_trajectory>
</example_3>
Prompts for PRDJudge in free development setting:
Please evaluate the {project_dir} project by running its tests and generating an evaluation report according to the evaluation criteria. The evaluation criteria are provided in evaluation/detailed_test_plan.json, and the project code is located in the src/ directory.
Please carefully read the code, rewrite the test file interface to adapt to the current code under src/, and offer reliable evaluation for current code.
The detailed evaluation report must be saved to reports/round2.jsonl in JSON format. Entries in the report should follow this structure:
"metric": "1.3 Menu Navigation - Export Results Submenu",
[…]
"score": 0,
"explanation": "When attempting to export results without having generated Huffman codes, the program does not enter the export submenu but instead prompts No available Huffman codes, please generate them first.’ and returns to the main menu, which does not meet the expected behavior."
"metric": "3.2 Unit Test - Generate Huffman Codes",
[…]
"score": 2,
"explanation": "The test command pytest src/tests/test_huffman.py::TestHuffman::test_generate_huffman_codes -v executed successfully, and the result was PASSED’, which matches the expected output Unit test passed’."
Please strictly follow the evaluation criteria in evaluation/detailed_test_plan.json, rewrite the inference in the evaluation files, run the relevant tests, and generate a comprehensive evaluation report as described above. Save the report to reports/round2.jsonl in the specified format.
Note: The prompt design for the Free Development mode requires PRDJudge to act as both an interface-adapter and an evaluator. As discussed in Section 5.4.4, this dual requirement introduces complexity and potential evaluation noise. We provide this prompt as a baseline for future research into unconstrained agent evaluation.
Appendix F Model Name Abbreviations
For brevity and formatting constraints—particularly within tables and figures—various foundation models and code agents are referred to by their abbreviated names throughout this paper. To prevent any ambiguity, we provide a comprehensive mapping between the official full names of the models and their corresponding abbreviations used in our text, tables, and figures.
Table 9 summarizes the nomenclature for all baseline models, backbone LLMs, and our evaluation agents.
Table 9. Full Name and Abbreviation Mapping for Models and Agents
Official Full Name Abbreviation(s)
Minimal Agent LLMs
Qwen3-Coder-480B-A3B Qwen3-Coder(480B), Qwen-Coder
Claude-4.5-Sonnet Claude-4.5, Claude
GPT-5.2 GPT-5.2
Gemini-3-Pro Gemini-3-Pro, Gemini
Kimi-K2-Instruct-0905 Kimi-K2
DeepSeek-v3.2 DeepSeek-V3.2
GLM-4.7 GLM-4.7
Minimax-M2 Minimax-M2
Commercial Agent LLMs
Gemini-2.5-Pro Gemini-2.5-Pro
GPT-5 GPT-5
Evaluation Agent LLM
Qwen3-Coder-30B-A3B Qwen3-Coder (30B), Qwen3-Coder 30B
Note: In Section 5.1, we explicitly declare the primary abbreviations for the Minimal Agents. However, in certain dense visualizations such as Figure 6 and Figure 7, further truncated names like “Qwen3”, “Claude”, and “Gemini” are employed to accommodate spatial limits.